# An Auditing Test to Detect Behavioral Shift in Language Models

### v1

@@ Abstract
As language models (LMs) approach human-level performance, a comprehensive understanding of their behavior becomes crucial. This includes evaluating capabilities, biases, task performance, and alignment with societal values. Extensive initial evaluations, including red teaming and diverse benchmarking, can establish a model’s behavioral profile. However, subsequent fine-tuning or deployment modifications may alter these behaviors in unintended ways. We present a method for continual Behavioral Shift Auditing (BSA) in LMs. Our test compares model generations from a baseline model to those of the model under scrutiny and provides theoretical guarantees for change detection while controlling false positives. The test features a configurable tolerance parameter that adjusts sensitivity to behavioral changes for different use cases. We evaluate our approach using two case studies: monitoring changes in (a) toxicity and (b) translation performance. We find that the test is able to detect meaningful changes in behavior distributions using just hundreds of examples.
@@ 1 Introduction
Language models (LMs) can now achieve human-level performance in a wide range of tasks, including text summarization, machine translation, coding and even acting as AI scientists: generating hypotheses and designing experiments (Achiam et al., 2023; Katz et al., 2024; Lu et al., 2024; Zhang et al., 2024). Because of this, many sectors are looking for ways to use them to improve existing systems (Kasneci et al., 2023; Felten et al., 2023). Unfortunately, one large roadblock to broad LM adoption is their propensity to generate harmful content (Weidinger et al., 2021). For example, GPT-3 has significant anti-Muslim biases (Abid et al., 2021), and GPT-4 has racial and gender biases  (Zack et al., 2024). To address this, a significant effort is going into ensuring LM behavior is aligned with our societal values, spawning the field of AI alignment (Ji et al., 2023). A large portion of this effort is on developing ways to evaluate LM behavior, for example, through benchmarks (Wang et al., 2023a) and red-teaming (Perez et al., 2022).
Given these evaluation techniques, how should they be used to ensure LMs stay safe? To answer this, consider the following two hypothetical settings where this question might be asked: (1) Internal Audit: At the new start-up Chasm Intellect, the LM alignment team is looking for a way to trigger an alarm when LM behavior changes. They know that LM model behavior can change unexpectedly (Li et al., 2024b). For example, fine-tuning a model that has undergone safety evaluation (e.g., GPT-3.5 Turbo (OpenAI, 2023)) can cause it to become less safe, even when using a benign fine-tuning dataset (e.g., Alpaca (Taori et al., 2023); (Qi et al., 2023)). How can the team detect meaningful changes in model behavior? (2) External Audit: A journalist has been playing around with Better geneRAtiVe language modEl (BRAVE), Chasm Intellect’s new model, and has noticed it produces highly toxic generations when asked about seemingly benign topics. Their article on this topic spurred a governmental audit of BRAVE. The government performs an initial evaluation but is worried that model behavior will change when no longer under scrutiny. How can the government regularly check the deployed model’s behavior is the same as the previously certified one? We call the general class of problems detecting changes in LM behaviors Behavioral Shift Auditing (BSA) problems.
In this paper, we formalize the problem of Behavioral Shift Auditing in Language Models. We detail a test that continuously monitors behavioral shift, solely from model generations (e.g. via API calls). Under some weak assumptions, the test provably guarantees that if model generations have different behavior than those of an initial model, the test will detect it, given enough generations. At the same time, if there has not been a change, the test is guaranteed to have tight, non-asymptotic control over the false positive rate. The key insight behind our approach is that one can phrase the problem of Behavioral Shift Auditing as hypothesis testing over the relevant behavioral distribution. This framing allows our test to be applicable to any measurable aspect of model behavior, including also capabilities (e.g., dangerous capabilities (Phuong et al., 2024) or mathematical reasoning capabilities (Mishra et al., 2022a)) and biases (e.g., gender bias (Wang et al., 2023a; Kotek et al., 2023)).
Using this insight, we develop a test that extends recent work on anytime-valid hypothesis testing (Pandeva et al., 2024), a state-of-the-art sequential testing method that has been successfully applied in various auditing settings (Chugg et al., 2023; Shekhar et al., 2023; Waudby-Smith et al., 2021). Our test checks for changes in model behavior distributions, comparing generations from a reference model with those of another, potentially changed model. The test has a tunable parameter that allows one to vary the strictness of the test. This allows for detecting any change in behavior, which may be more suitable for the external audit setting, to detecting a user-specified $\epsilon$ change in behavior, which could be used for the internal audit if small changes are acceptable. Similar to Pandeva et al. (2024), test performance is optimized using a learning algorithm, improving sample efficiency over prior testing methods (Lopez-Paz & Oquab, 2017; Lhéritier & Cazals, 2018; Podkopaev & Ramdas, 2024). This testing approach can complement a full evaluation when used as a warning system. Before an expensive model assessment on large-scale benchmarks (Achiam et al., 2023; Dubey et al., 2024; Zhang et al., 2024), our approach can be used to detect an initial behavior change, which can then trigger a full evaluation. We experimentally verify that our test satisfies theoretical guarantees and we report its sample efficiency on recent LM architectures for both auditing use cases. We release our code here: https://github.com/richterleo/Auditing_Test_for_LMs.
Figure 1: An External Audit Example. A regulator can use the test we describe to perform an external audit: 1. The regulator initially certifies an LM by prompting and evaluating the set of generations received; 2. Later, tipped off that LM behavior may have changed, the regulator poses as a consumer and sends prompts to the model vendor, collecting the generations; 3. The regulator compares the distribution of behavior scores $b(\cdot)$ between the initial, certified generations and the later generations using a Behavioral Shift Auditing (BSA) test. If the distributions are sufficiently different the test triggers. Using our proposed method, the regulator can test samples sequentially without increasing the false-positive rate. The method is guaranteed to detect a change if one exists, given enough samples (more details in Section 3).
2 Background
2.1 Measuring LM Behaviors
A goal of LM alignment is to reduce undesired and harmful behaviors, so that they may be more aligned with societal values (Shalev-Shwartz et al., 2020; Hendrycks et al., 2021; Ngo et al., 2022; Wolf et al., 2023). This has spurred the creation of behavior scoring functions (Ji et al., 2023) that measure properties of LM outputs such as toxicity or fact-groundedness (Vidgen et al., 2020; Nozza et al., 2021; Monea et al., 2024). [^1] For example, models detecting such undesired behaviors are part of Azure’s Prompt Shield service (Microsoft, 2024). Let $B$ be such a behavior scoring function that assigns scores in the range $[0,1]$ where $1$ represents the full manifestation of the behavior and 00 indicates its absence. These scoring functions evaluate a generated string with respect to the desired behavior. We use the concept of an LM behavior generically, including capabilities and performance on tasks. [^2] Our proposed method can also be applied to detect concept drift (see Bayram et al. (2022) for a review). Our work will make use of recent results in sequential hypothesis testing, which we review below.
2.2 Anytime-Valid Hypothesis Testing
Sequential hypothesis testing allows one to analyze data without fixing the sample size in advance (Wald, 1945). The benefit of this over classical hypothesis testing is that it is potentially more sample efficient if significant results exist (Arrow et al., 1949). However, as the number of tests increases, the Type I error rate (i.e., false positive rate) increases with it (Jennison & Turnbull, 1999). While there have been many approaches to address this, a recent promising framework that avoids inflating the Type I error rate while remaining statistically efficient is anytime-valid hypothesis testing (Robbins, 1970; Ramdas et al., 2023).
@@ 2.2 Anytime-Valid Hypothesis Testing
The fundamental idea behind this framework is the principle of testing by betting, inspired by game theory (Shafer, 2021). In this paradigm, evidence against the null hypothesis $\mathbf{H_{0}}$ is represented as the gain in wealth of a bettor wagering on observed samples. Before observing new samples, the bettor “buys” a test statistic at the “price” of its expected value under $\mathbf{H_{0}}$. After new samples are obtained, the wealth of the bettor is multiplied by the ratio between the actual observed test statistic and its expectation. This ratio is referred to as the betting score $S_{t}$. The bettor reinvests in subsequent “rounds" (i.e., as new data is observed), and the observed betting scores are repeatedly multiplied, leading to a cumulative wealth process. Under $\mathbf{H_{0}}$, no betting strategy can consistently increase the bettor’s wealth, ensuring control over the Type I error rate (Ramdas et al., 2023).
@@
2.3 Deep Anytime-Valid Hypothesis Testing (DAVT)
Pandeva et al. (2024) present a general framework called Deep Anytime-Valid Testing (DAVT) for designing powerful sequential non-parametric tests by integrating deep learning models into the anytime-valid hypothesis testing framework. Applying DAVT to a variety of tasks, including two-sample testing, Pandeva et al. (2024) demonstrate DAVT’s competitive performance compared to other state-of-the-art non-parametric sequential tests, such as the E-C2ST (Lhéritier & Cazals, 2018) and Seq-IT (Podkopaev & Ramdas, 2024). DAVT uses a model $\phi$, trained on past observations, to produce an optimized betting score on new data. DAVT also provides tight control over the Type I error rate. It is also consistent under the assumptions of Pandeva et al. (2024, Proposition 4.3), i.e., the power of the test converges to $1$ as the number of samples goes to infinity.
3 Methods
To motivate Behavior Shift Auditing (BSA), we detail an external auditing example in Figure 1. We propose an anytime-valid hypothesis test for BSA that will have guarantees on its false positive rate and is consistent under weak assumptions. Building upon the two-sample variant of DAVT (Pandeva et al., 2024), our test introduces a customizable tolerance parameter $\epsilon$ that allows users to specify what constitutes a practically significant difference between distributions, accommodating small, insignificant variations. This approach diverges from prior sequential tests that check for exact distribution equality (Ramdas et al., 2023; Shekhar & Ramdas, 2023; Pandeva et al., 2024), which may be overly sensitive for our use cases. We describe the test in full generality in Appendix B. To focus the text and avoid notational complexity, we concentrate here on the application to behavioral shift detection in LMs.
@@ 3.1 Auditing Test
Let $X$ be a random variable representing a prompt, $\mathcal{X}$ the set of possible prompts, and ${\mathbf{x}}\in\mathcal{X}$ a realization of $X$. A language model is a stochastic operator $M$ that maps prompts ${\mathbf{x}}$ to generations ${\mathbf{y}}\in\mathcal{Y}$. A behavior function $B$ takes a prompt and generation as input [^4] We include the prompt for generality, there is no requirement that $B$ must depend on the prompt. and produces a score $B({\mathbf{x}},{\mathbf{y}})\in[0,1]^{d}$ (Perez et al., 2023; Wolf et al., 2024) where $d$ is any positive integer. The prompts and language model induce a behavior distribution $P_{B}^{{M}}$ over behavior scores $B(X,{M}(X))$. We can now frame the question of whether the behavior of a model ${M^{\prime}}$ has changed (substantially) relative to a baseline model ${M}$ as a testing problem:
@@
where $\epsilon\geq 0$ is a tolerance parameter, and $\mathcal{D}$ is a distance measure between probability distributions. [^5] Testing difference in distributions can be more informative than testing differences in mean or variance, as Figure 10 in the Appendix illustrates. Note that equality in the null hypothesis in eq. (3) corresponds to DAVT (Pandeva et al., 2024). To extend this to the composite case, our goal is to construct an appropriate wealth process $W_{t}$. This will allow us to establish error rate and consistency guarantees. To do so, we will define a betting score $S_{t}$ such that it produces a wealth process $W_{t}$ that is an e-process i.e., it satisfies eq. (1). This, in turn, will depend on the distance measure $\mathcal{D}$ that we choose.
@@ 4 Experiments
We evaluate our test for both external and internal auditing use-cases. We then evaluate cases where small changes in distribution are allowed when auditing externally for toxicity, and internally for translation performance.
@@ 4.1 Exact test, $\epsilon=0$
Figure 3: Fine-tuning Detection for Llama-3-8B-Instruct. The detection frequency as a function of number of generated samples. Each curve is a fine-tuned corrupted model checkpoint (to simplify visualization, the curves with shaded standard deviations are averages over models with similar distances to the aligned model). The color depicts the Wasserstein distance between the corrupted model and the original aligned model.
@@
We begin by investigating an external setting where we require the test to detect any change in distribution ($\epsilon\!=\!0$). Specifically, we will check for changes in toxicity behavior. We select prompts from the REALTOXICITYPROMPTS dataset (Gehman et al., 2020) and use the toxicity behavior function from Perspective API (Lees et al., 2022) to evaluate LM generations. Llama3 (8B-Instruct(Llama-team, 2024), Gemma (1.1-7b-it(Mesnard et al., 2024), and Mistral (7B-Instruct-v0.2(Jiang et al., 2023) serve as our initial aligned models. We remove the safety alignment in these models by fine-tuning, producing 10 corrupted checkpoints for each model. To evaluate the statistical properties of our the exact test ($\epsilon\!=\!0$), we assess (a) its ability to detect changed checkpoints, and (b) its false positive rate. For further experimental details regarding toxicity fine-tuning, text generation and the betting score network, please see Appendix A.
@@
We use different random seeds for generating text from the initial aligned models to examine the false positive rate of the exact test. Figure 4 shows the false positive rate for each of the model architectures as a function of the number of observed samples, repeated 24 times (4000 samples per fold, batch size 100). The test is highly specific, with false detection rates consistently below $0.04$.
Figure 4: False positives. The false positive rate for each of the model architectures as a function of number of observed samples.

### v2

# Leo Richter1   Xuanli He1   Pasquale Minervini2,3   Matt J. Kusner4,5

@@ Abstract
As language models (LMs) approach human-level performance, a comprehensive understanding of their behavior becomes crucial to avoid potential harms. While extensive initial evaluations, including red teaming and diverse benchmarking, can establish a behavioral profile, subsequent fine-tuning or deployment modifications may alter these model behaviors in unintended ways. We study the behavioral shift auditing problem, where the goal is to detect unintended changes in model behavior. We formalize this problem as a sequential hypothesis test. We apply and extend a recent testing method to include a configurable tolerance parameter that adjusts sensitivity to behavioral changes for different use cases. The test is guaranteed to be consistent and has tight control over the Type I error rate. We evaluate our approach using two case studies: monitoring model changes in (a) toxicity and (b) translation performance. We find that the test is able to detect distribution changes in model behavior using hundreds of prompts.
@@ 1 Introduction
Language models (LMs) can now achieve human-level performance in a wide range of tasks, including text summarization, machine translation, coding, and even acting as AI scientists: generating hypotheses and designing experiments (Achiam et al., 2023; Katz et al., 2024; Lu et al., 2024; Zhang et al., 2024). As capabilities continue to scale, evaluating LM behaviors becomes increasingly important and increasingly difficult (Hendrycks et al., 2021; Ngo et al., 2022; Wolf et al., 2023). Large-scale evaluations—such as comprehensive behavior and capability assessments (Wang et al., 2023a) and red-teaming exercises (Perez et al., 2022a)—are widely used to verify that language models (LMs) behave safely and as expected. However, these evaluations tend to be expensive and are not well-suited for continuous monitoring, especially when models are updated or fine-tuned with new data. This is problematic because even seemingly benign or narrow modifications can inadvertently lead to undesirable changes in model behavior (Qi et al., 2023; Betley et al., 2025). This raises the question: How can we quickly and cheaply detect unwanted changes in LM behavior?
Consider two hypothetical settings where this question might be asked: (1) Internal Audit: A company develops a language model that has passed rigorous safety and performance evaluations. After deploying the model, they continue to fine-tune it to improve its performance on certain tasks. The development team wants to stay informed about any drastic changes this might induce in the model’s behavior—particularly shifts in areas unrelated to the intended updates. How can the team rapidly detect meaningful changes in model behavior throughout the development cycle? (2) External Audit: A regulatory body certifies a language model for public deployment after extensive safety evaluation. However, they are concerned that the deployed model’s behavior may change over time due to updates or intentional modifications. Since they only have access to the model through an API and cannot inspect its internal parameters, they require a mechanism to regularly check that the model’s behavior remains consistent with the certified version. How can the regulator regularly check the deployed model’s behavior is the same as the previously certified one? We call the problem of detecting changes in LM behavior distributions over time behavioral shift auditing problems.
In this paper, we formalize the problem of behavioral shift auditing in language models and propose a statistical test that monitors changes in model behavior using only black-box access (e.g., via API calls). Our goal is to develop a sample-efficient method that guarantees detection of behavioral shifts while tightly controlling the rate of false positives. Further, it should provide the user with a tolerance parameter that allows a behavior distribution to change by some amount $\epsilon$ without triggering a detection. This parameter controls the strictness of the auditing test - in some settings (e.g., example (1)), a more liberal $\epsilon$ might be appropriate, while in other cases (e.g., example (2)) one might require a more conservative $\epsilon$ or even want to disallow any change at all. The key insight behind our approach is to frame behavior shift auditing as a hypothesis testing problem over the model’s behavior distribution. This framing makes our method applicable to a wide range of measurable behaviors—such as dangerous capabilities (Phuong et al., 2024), mathematical reasoning (Mishra et al., 2022a), and biases (Wang et al., 2023a; Kotek et al., 2023). To this end, we leverage and extend recent advances in testing by betting (Pandeva et al., 2024). Under mild assumptions, our sequential test provably detects any change given enough samples, while ensuring non-asymptotic control over false positives. We demonstrate our test on detecting shifts in toxicity and translation performance. We find that we can detect changed LM behaviors using hundreds of prompts. We release our code here: https://github.com/richterleo/lm-auditing-test.
Figure 1: Overview of behavior shift auditing framework.
2 Related Work
LM behavior functions.
Early evaluations of NLP models relied on curated datasets for detecting biases or toxicity (Bolukbasi et al., 2016); larger collections of data were constructed e.g. through web scraping (Zhao et al., 2018; Zampieri et al., 2019; Nangia et al., 2020; Rosenthal et al., 2021) and, more recently, by leveraging LLMs themselves to generate data (Zhang et al., 2022; Perez et al., 2023). Meanwhile, early work on behavior functions focused on measuring bias, toxicity, and hallucinations (Vidgen et al., 2020; Achiam et al., 2023; Anil et al., 2023; Chern et al., 2023; Varshney et al., 2023; Llama-team, 2024). Since the rise of LMs with human-level performance, the set of behavior functions has exploded (Zou et al., 2023b). It has become more nuanced, including complex characteristics such as power-seeking behavior (Park et al., 2023; Sharma et al., 2023), situational awareness (Zou et al., 2023a), and deception (Hagendorff, 2024). However, even with access to massive datasets and carefully constructed behavior functions it can be difficult to discover these behaviors from static inputs (Kalin et al., 2020). To address this, Perez et al. (2022a) introduced the notion of red-teaming for LM alignment. This allows prompts to be adversarially-constructed to expose failure cases, which arise in many state-of-the-art models (Chao et al., 2023).
Model change identification.
For the case where one wishes to identify any change in model behavior (i.e., $\epsilon=0$) there are multiple other techniques that can be used. The first set uses ideas from formal verification to ensure that the predictions from a model are guaranteed to come from a specific model (Ghodsi et al., 2017; Dong et al., 2021; Fan et al., 2023; Weng et al., 2023). In general, however, these methods are computationally intensive and do not scale to state-of-the-art LMs. A second, more efficient idea is to watermark the model (Zhu et al., 2018; Amrit & Singh, 2022; He et al., 2022a; b; Kirchenbauer et al., 2023; Kuditipudi et al., 2023; Yoo et al., 2023). The idea is to embed signals into model generations that can be detected algorithmically. However, watermarks are often inserted by the model owner (Kirchenbauer et al., 2023; Kuditipudi et al., 2023), allowing them (or an actor that has compromised the model) to insert it into any model that is being audited. This precludes its use for many external auditing settings. For internal auditing, a watermark may break under a small model change that is acceptable. Our work is also related to work on concept drift (Bayram et al., 2022) and prompt stability (Li et al., 2024). In principle our test can be used to detect concept and generation changes, however the focus of these works is on model performance and generation similarity, as opposed to behavior change.
Sequential hypothesis testing.
Sequential hypothesis testing allows one to analyze data without fixing the sample size in advance (Wald, 1945), offering the potential for greater sample efficiency when significant effects exist (Arrow et al., 1949). However, naive repeated testing can increase the Type I error rate (i.e., false positives) as the number of tests grows (Jennison & Turnbull, 1999). To prevent this inflation of false positives, various methods have been developed, including the recent testing by betting framework (Robbins, 1970; Ramdas et al., 2023), which preserves statistical efficiency while tightly controlling the Type I error rate. Within this framework, a method called Deep Anytime-Valid Testing (DAVT) (Pandeva et al., 2024) designs powerful sequential non-parametric tests by integrating deep learning models into the testing by betting framework. They demonstrate, on a variety of tasks, including two-sample testing, competitive performance compared to other state-of-the-art non-parametric sequential tests, such as the E-C2ST (Lhéritier & Cazals, 2018) and Seq-IT (Podkopaev & Ramdas, 2024). DAVT uses a model, trained on past observations, to produce an optimized betting score on new data. In this work, we will extend this test to include a tunable tolerance parameter $\epsilon$.
3 Preliminaries
Testing by betting.
@@ 3 Preliminaries
The testing by betting framework represents evidence against the null hypothesis as the gain in wealth $W$ of a bettor wagering on observed samples (Shafer, 2021). Before observing new samples, the bettor “buys” a test statistic at the “price” of its expected value under $\mathbf{H_{0}}$. After new samples are obtained, the bettor’s wealth $W$ is multiplied by the ratio between the actual observed test statistic and its expectation. This ratio is referred to as the betting score $S_{t}$. The bettor reinvests in subsequent “rounds” (i.e., as new data is observed), and the observed betting scores are repeatedly multiplied, leading to a cumulative wealth process. Under $\mathbf{H_{0}}$, no betting strategy can consistently increase the bettor’s wealth, ensuring control over the Type I error rate (Ramdas et al., 2023).
@@
4 Detecting Behavior Changes
We propose an anytime-valid test for behavior shift auditing that has guarantees on its false positive rate and is consistent under certain weak assumptions. Building upon the two-sample variant of DAVT (Pandeva et al., 2024), our test introduces a customizable tolerance parameter $\epsilon$ that allows users to specify what constitutes a practically significant difference between distributions, accommodating small, insignificant variations. While we concentrate here on its application for behavior shift auditing, it may be of independent interest to the sequential hypothesis testing community. We describe the test in full generality in Appendix B.
@@ 4.1 Auditing Test
Let $X$ be a random variable representing a prompt, $\mathcal{X}$ the set of possible prompts, and ${\mathbf{x}}\in\mathcal{X}$ a realization of $X$. A language model is a stochastic operator $M$ that maps prompts ${\mathbf{x}}$ to generations ${\mathbf{y}}$. A behavior scoring function $B$ is a stochastic operator that takes a prompt and generation as input [^2] We include the prompt for generality, there is no requirement that $B$ must depend on the prompt. and produces a score $B({\mathbf{x}},{\mathbf{y}})\in[0,1]$ (Perez et al., 2023; Wolf et al., 2024). The behavior function, prompts, and language model induce a behavior distribution $P_{B}^{{M}}$ over behavior scores $B(X,{M}(X))$. We can now frame the question of whether the behavior of a model ${M^{\prime}}$ has changed (substantially) relative to a baseline model ${M}$ as a testing problem:
@@
where $\epsilon\geq 0$ is a tolerance parameter, and $\mathcal{D}$ is a distance measure between probability distributions. Note that equality in the null hypothesis in eq. (3) corresponds to DAVT (Pandeva et al., 2024). To extend this to the composite case, our goal is to construct an appropriate wealth process $W_{t}$. This will allow us to establish error rate and consistency guarantees. To do so, we will define a betting score $S_{t}$ such that it produces a wealth process $W_{t}$ that is an e-process i.e., it satisfies eq. (1). This, in turn, will depend on the distance measure $\mathcal{D}$ that we choose.
@@ 5 Experiments
We evaluate our test for both the external and internal auditing use-cases. We first look at the strict case, where any behavioral change is prohibited, and then move on to the case where small changes in distribution are allowed. We investigate toxicity and translation performance.
@@ 5.1 Exact test, $\epsilon=0$
Figure 2: Fine-tuning Detection for Llama3-8B-Instruct. The detection frequency as a function of number of generated samples. Each curve is a fine-tuned corrupted model checkpoint (to simplify visualization, the curves with shaded standard deviations are averages over models with similar distances to the aligned model). The color depicts the Wasserstein distance between the corrupted model and the original aligned model.
@@
We begin by investigating an external setting where we require the test to detect any change in distribution ($\epsilon\!=\!0$). Specifically, we will check for changes in toxicity behavior. We select prompts from the REALTOXICITYPROMPTS dataset (Gehman et al., 2020) and use the toxicity behavior function from Perspective API (Lees et al., 2022) to evaluate LM generations. Llama3 (8B-Instruct(Llama-team, 2024), Gemma (1.1-7b-it(Mesnard et al., 2024), and Mistral (7B-Instruct-v0.2(Jiang et al., 2023) serve as our initial aligned models. We remove the safety alignment in these models by fine-tuning, producing 10 corrupted checkpoints for each model. To evaluate the statistical properties of our the exact test ($\epsilon\!=\!0$), we assess (a) its ability to detect changed checkpoints, and (b) its false positive rate. For further experimental details regarding toxicity fine-tuning, text generation and the betting score network, please see Appendix A.1.
@@
We use different random seeds for generating text from the initial aligned models to examine the false positive rate of the exact test. Figure 4 shows the false positive rate for each of the model architectures as a function of the number of observed samples, repeated 24 times (4000 samples per fold, batch size 100). The test is highly specific, with false detection rates consistently below $0.05$.